if 2 person can paint 2 house in 2 hours then, then how many person can paint 18 house in 6 hours
Answer by ankor@dixie-net.com(22740)

if 2 person can paint 2 house in 2 hours
Find how many man-hrs per house
2 * 2 / 2 = 2 man hrs per house
:
then, then how many person can paint 18 house in 6 hours
let m = no. men to accomplish this
6m = 18(2)
6m = 36
m = 36/6
m = 6 men